OpenStudy (anonymous):

.A body traveling along a straight line traversed one third of the total distance with a velocity 4m/s the remaining part of the distance was covered with a velocity 2m/s for half time and with velocity 6m/s for the other half of timethe mean velocity average over the whole time of motion is

OpenStudy (mos1635):

 4*t1=1/3 l(el) 2*t2+6*t2=2/3 l=>8t2=2/3l from that calculate t1+t2 and aplie that to mean velocity........